OpenStudy (anonymous):

simplify the trig identity. 1-((cos^2 theta)/(1+sintheta))=sin theta

OpenStudy (isaiah.feynman):

Sure the question is correct? @ashmibri

OpenStudy (anonymous):

yes i'm sure

OpenStudy (isaiah.feynman):

Please just check one more time.

OpenStudy (anonymous):

\[1- (\cos ^{2}\theta \div 1+\sin \theta) = \sin \theta \]

OpenStudy (isaiah.feynman):

PROVED IT!

OpenStudy (isaiah.feynman):

\[1-\frac{ \cos^{2}\theta }{ 1+\sin \theta}\]

OpenStudy (isaiah.feynman):

\[\frac{ 1+\sin \theta-\cos^{2}\theta }{ 1+\sin \theta }=\frac{ 1+\sin \theta-1+\sin^{2}\theta }{1+\sin \theta }=\frac{ \sin \theta+\sin^{2}\theta}{ 1+\sin \theta }=\sin \theta \]

OpenStudy (anonymous):

It actually comes out really clean. See Picture

OpenStudy (anonymous):

thank you very much

OpenStudy (isaiah.feynman):

My medal?

OpenStudy (anonymous):

Medal for what??